Lee has 3 kids the oldest is twice as old as the youngest the middle child is 5 years older than the youngest the sum is 57
Answer by tenkun(12)

let age of yougest = x
the age of middle child = x+5
the age of oldest child = 2*x
given that sum of their age = 57
so x + (x+5) + 2*x = 57
4*x + 5 =57
4*x = 52
x = 52/4
x= 13
hence age of youngest child = 13
age of middle child = 13 + 5 =18
age of oldest child = 2*13 =26